# Data-driven Energy Efficiency Modelling in Large-scale Networks: An Expert Knowledge and ML-based Approach

### v1

@@ I Introduction
The potential of these network energy saving solutions is, however, suboptimally exploited in current fourth generation (4G) and 5G network deployments due to the complexity of their optimization. The modelling and further optimization of network energy saving solutions are complex problems, largely unsolved in both industry and academia. The main challenges emanate from (i) the large number of cells and parameters per cell to configure, (ii) the stochastic and non-stationary nature of end-user traffic demands and the wireless channel, and (iii) the intricate coupling/trade-offs between energy consumption and network/UE performance (in terms of coverage quality, throughput, reliability and latency) [6].
@@
Various methodologies have been employed to tackle the issue of optimizing network energy saving solutions. However, none of these techniques have completely overcome the challenges at hand.
@@
Currently, expert knowledge still remains a primary means for optimizing network parameters. Experts rely on cell-level statistics [7], gathered from BSs, as well as call detail records [8] and drive tests [9] measurements to evaluate network performance and UE quality of service at various locations, and take decisions. However, CDRs and DTs necessitate large measurement campaigns and specialized equipment, which make them economically expensive and diminish their worth. To alleviate this cost challenge, the 3GPP introduced minimization of DT [10], which employs UE geolocated radio measurements to inexpensively assess network performance. However, only a few UEs enable this feature nowadays, and they usually do it for a short period of time. Privacy is also a concern. Nevertheless, it is essential to acknowledge that cell-level statistics, CDRs, DTs, and minimization of DT data alone are not sufficient for network optimization. These metrics assess network performance only for the specific configuration —or, at most, a few configurations— set during the measurement campaigns. They do not provide insights into network performance across the billions of other potential network configurations. Consequently, state of the art trial and error optimization approaches based on expert knowledge as well as cell-level statistics, CDRs and DT-based data cannot accommodate for the requirements of today’s real world 4G/5G large-scale network optimization.
@@
System-level simulators are another widely used option for assisting network optimization [11]. Precise modeling of the propagation environment and protocol stack are crucial to obtain an accurate evaluation and generalization of network performance to any network configuration. Available tools for modeling the wireless propagation environment range from basic models that solely rely on statistics to advanced ray-tracing systems [12]. However, these precise tools often require hard-to-obtain map information and time-consuming operations, making them expensive. Examples of such radio propagation prediction tools are [13, 14]. With regard to the protocol stack, frameworks such as NS-3 5G-LENA [15] and its evolution [16] have gained popularity due to their capabilities. However, they do not model the complexity of a true product. Importantly, they do not take into account the heterogeneity —and different performance— of off-the-shelf BS and UE products.
Theoretical tools have also been developed to aid network energy efficiency optimization. Importantly, optimal shutdown policies for a single server with bursty traffic were derived in [17]. The optimal shutdown policy was shown to be a two-threshold policy, with one threshold to drive the sleep procedure and a different one to manage the cell wake up. Even if there is no theoretical results available for more than one server, the findings of this work are in line with today´s carrier shutdown solutions, which implement two different sets of conditions for shutdown and reactivation (see Section II-B). When theoretically studying more complex multi-cell macrocellular networks, we should highlight the coverage, capacity and energy efficiency trade-offs derived in [18], [19] and [20] for large-scale cellular networks with massive multiple-input multiple-output (mMIMO), carrier aggregation (CA) and joint mMIMO and CA capabilities, respectively. In [21], the associated energy efficiency trade-offs of a macrocellular network equipped with shutdown capabilities and a small cell network were compared, and the pros and cons of these two different deployments discussed. The further analysis in [22] showed the benefits of deploying advanced sleep modes at the small cells, proposing multiple carrier shutdown strategies, and comparing them in terms of complexity, blocking rate probability, throughput and energy efficiency. From an optimization perspective, the authors of [23] investigated how to fine-tune network energy efficiency by jointly managing the long-term cell activation, UE association and power control in a heterogeneous network with mMIMO capabilities, and proposed a distributed solution based on game theory, proven to converge to the Nash equilibrium. The work in [24] also modelled hardware failure rate due to cell (de)activation, and proposed a heuristic to control the BS shutdown, which minimises the acceleration over time of the hardware failure rate, while satisfying the UE quality of service demands. Despite the value of the models generated and the conclusions drawn in all this research, a considerable number of assumptions were necessary due to the theoretical nature and the complexity of large-scale networks. For example, the carrier shutdown implementations considered in these studies deviate significantly from the practical solutions deployed in the field. Consequently, these modeling and optimization frameworks are not well-suited for the real-world optimization of practical network energy-saving solutions.
In recent years, the 3GPP has initiated efforts to tackle the aforementioned issues by working on enabling a new data-driven network modeling and optimization paradigm [25]. This involves identifying the necessary measurements and protocols for the development of advanced data-driven network optimization approaches [26], and leveraging ML to predict network performance and optimize various aspects, including energy efficiency [27]. Despite these strides, efficient realization and validation of this paradigm require careful consideration of design aspects, standardization, and continued innovation. To date, the research efforts on this 3GPP tools have mainly focused on developing ML solutions that leverage data generated through numerical simulations and/or test beds to design new physical layer mechanisms [28] and optimize upper layer functionalities in cellular networks [29]. In contrast, only few studies have investigated real network data for modelling and optimizing mobile communication systems. In this context, Xing et al. have proposed a neural network framework to model the downlink (DL) spectral efficiency of 5G mMIMO BSs based on reference signal received power (RSRP) collected from DTs data [30]. This work is one of the first to statistically characterize network performance by solely using this type of data. However, spectral efficiency alone cannot capture the impact of energy saving strategies on network performance as it does not consider, among others, the amount of bandwidth used per cell and allocated per UE. There is a need for an end-to-end framework that fully models the trade-offs between energy consumption and other network key performance indicators, such as coverage, network load and cell/UE throughput, and considers complex interactions amongst energy saving solutions, mobility management and others.
@@
In this paper, to help addressing the network energy saving solution optimization challenge, we propose a novel data-driven network energy efficiency modelling framework, referred to as SRCON [31], which uses measurement data from live networks to jointly fit (i) ML-based, black-box and (ii) expert-based, white-box models222 In contrast to a black-box model, in a white-box model logic, functioning and programming steps are transparent and known. As a result, it’s decision making process is interpretable. that can
@@
  • generalise the modelling of network energy efficiency and UE quality of service to any possible network energy saving solution configuration, even those not observed in the training data.
@@
With respect to the available literature, the novelty of this paper resides in the following aspects:
@@
  • Novel Modelling Approach: Unlike existing literature, this work uniquely focuses on modeling network energy consumption and UE data rates within the context of a practical carrier shutdown solution in a real and extensive heterogeneous network, encompassing hundreds of 4G and 5G cells. No expensive drive testing and/or ray tracing-based simulation tools are required.
@@
  • Decomposed Modelling Framework: This paper introduces a novel approach by decomposing the overall modeling problem into distinct ML- and expert-based modeling subproblems. Outputs from the expert-based models serve as inputs for ML-based models. This careful decomposition enables accurate generalization of network energy consumption and UE performance predictions to any configuration of the carrier shutdown solution, even for scenarios not encountered in the data.
  • Data and Model Construction Pipeline: A comprehensive description is provided for the data required and the pipeline employed to construct ML-based models for network energy consumption and UE data rates. This approach is essential for capturing the diverse performance characteristics of various network products and end-user handsets existing in a network. The different problems require different methodologies.
  • Innovative Expert-Based ABM Model: The introduction of an innovative agent-based model (ABM) represents a significant contribution. This expert-based model effectively captures the stochasticity of the network and intricately models the behavior of carrier shutdown in a wide-area network, with an affordable complexity. The ABM enables accurate predictions, leveraging well-established telecom theory, and provides inputs to ML-based models for any carrier shutdown solution configuration.
@@
  • Data-Driven Comparative Results: The paper concludes with data-driven results that highlight the substantial benefits of the proposed modeling approach. Comparative analysis against a benchmark modeling tool, employed by a top-tier network operator for optimizing carrier shutdown solutions, showcases the superior performance and efficacy of the presented models.
@@ II-B Reference carrier shutdown (de)activation logic
In the reference implementation used in this paper, and assuming that a capacity cell, $c$, is paired with only one coverage cell, $b(c)$, carrier shutdown can be autonomously activated in capacity cell, $c$, when
@@ III Network Energy Efficiency Modelling
As can be inferred from previous sections, the network energy efficiency modelling problem deals with multiple objectives (e.g. energy consumption, network coverage, network/UE rate) and optimization variables (e.g. carrier shutdown and A4 handover parameters), and is particularly challenging due to its large-scale, stochasticity and non-stationarity as well as the complex coupling between cells and the intricate trade-offs between energy consumption and network/UE performance. For instance, the shutdown of a cell may save energy, but impacts, not only the performance of the UEs connected to such cell, but also the overall network coverage and the performance of those UEs connected to the neighboring cells as well as the possibility of nearby capacity cells to shutdown. In practice, these challenges make the utilization of precise networking models essential to perform a rigorous network energy efficiency optimization.
@@
To address these challenges, SRCON —our proposed modelling approach— combines a variety of wireless as well as data-driven and ML concepts. Rather than attempting to replicate network behaviours at a time-scale of milliseconds, which is fundamentally infeasible, considering the aforementioned modelling challenges in practical networks, SRCON emulates network behaviours in a statistically indistinguishable manner. This statistical indistinguishability lays the ground for a practical and accurate overall assessment of network/UE performance, allowing the use of sufficient statistics —instead of all available samples— through an efficient processing of massive data sets. To enable an accurate generalization, and deal with the heterogeneity and complexity of devices, SRCON uses measurement data from live networks to jointly fit ML-based, black-box and expert-based, white-box models, as hinted earlier. For our network energy efficiency modelling problem, in more details, we propose to model BS energy consumption and UE rate statistics using customised ML-based modeling approaches to the available data to capture the particularities of the specific BS products and off-the-shelf UEs in the area of study (e.g. not all BS product versions have the same energy consumption characteristics, not all UEs have the same decoding capabilities). A systematic feature importance analysis was used to identify the most relevant input features of such two ML-based models, e.g. the amount of time for which capacity cells will be in carrier shutdown, the number of UEs transferred to each neighbouring cell when carrier shutdown takes place, the resulting PRB load in those neighbouring cells. To allow generalisation, a customised expert-based model is used to derive the inputs of the ML-based models for any carrier shutdown and A4 handover parameter configuration. In other words, said white-box model takes as input, among others, the carrier shutdown and A4 handover parameter configuration, and provides as output the inputs of our black-box models. In this manner, we can drive universal ML-based BS energy consumption and UE rate predictions, and further derive network energy efficiency.

### v2

@@ I Introduction
The potential of these network energy saving solutions is, however, suboptimally exploited in current fourth generation (4G) and 5G network deployments due to the complexity of their optimization. The modelling and further optimization of network energy saving solutions are intricate problems, largely unsolved in both industry and academia. The main challenges emanate from (i) the large number of cells and parameters per cell to configure, (ii) the stochastic and non-stationary nature of end-user traffic demands and the wireless channel, and (iii) the intricate coupling/trade-offs between energy consumption and network/UE performance (in terms of coverage quality, throughput, reliability and latency) [6].
Effectively optimizing network energy saving solutions, not only enhances network energy efficiency and reduces operator costs, but also significantly contributes to environmental sustainability.
I-A Literature review
@@
Various methodologies have been employed to tackle the issue of optimizing network energy saving solutions. None of these techniques, however, have completely overcome the challenges at hand. In the following, we review the most common approaches to network energy saving modelling and optimization, introducing the role of ML.
I-A1 Expert knowledge
@@
Currently, expert knowledge still remains a primary means for optimizing network parameters. Experts rely on cell-level statistics [7], gathered from BSs, as well as call detail records [8] and drive tests [9] measurements to evaluate network performance and UE quality of service at various locations, and take decisions. However, CDRs and DTs necessitate large measurement campaigns and specialized equipment, which make them economically expensive and diminish their worth.
To alleviate this cost challenge, the 3GPP introduced minimization of DT [10], which employs UE geolocated radio measurements to inexpensively assess network performance. However, only a few UEs enable this feature nowadays due to privacy concerns, and they usually do it for a short period of time.
Nevertheless, it is essential to acknowledge that cell-level statistics, CDRs, DTs, and minimization of DT alone are not sufficient for network optimization. These metrics assess network performance only for the specific configuration —or, at most, a few configurations— set during the measurement campaigns. They do not provide insights into network performance across the billions of other potential network configurations. Consequently, state of the art trial and error optimization approaches based on expert knowledge as well as cell-level statistics, CDRs and DT-based data cannot accommodate for the requirements of today’s real world 4G/5G large-scale network optimization.
I-A2 System-level simulation
System-level simulation is another widely used option for assisting network optimization [11]. Precise modeling of the propagation environment and protocol stack are crucial to obtain an accurate evaluation and generalization of network performance to any network configuration.
@@ I-A2 System-level simulation
Available tools for modeling the wireless propagation environment range from basic models that solely rely on statistics to advanced ray-tracing systems [12]. However, these precise tools often require hard-to-obtain map information and time-consuming operations, making them expensive. Examples of such radio propagation prediction tools are [13, 14].
With regard to the protocol stack, frameworks such as NS-3 5G-LENA [15] and its evolution [16] have gained popularity due to their capabilities. Nevertheless, they fail at modelling the complexity of true products. For example, they do not take into account the heterogeneity —and different performance— of off-the-shelf BS and UE products.
I-A3 Theoretical tools
Theoretical tools have also been developed to aid network energy efficiency optimization. Importantly, optimal shutdown policies for a single server with bursty traffic were derived in [17]. The optimal shutdown policy was shown to be a two-threshold policy, with one threshold to drive the sleep procedure and a different one to manage the cell wake up. Even if there is no theoretical results available for more than one server, the findings of this work are in line with today´s carrier shutdown solutions, which implement two different sets of conditions for shutdown and reactivation (see Section II-B).
When theoretically studying more complex multi-cell macrocellular networks, we should highlight the coverage, capacity and energy efficiency trade-offs derived in [18], [19] and [20] for large-scale cellular networks with massive multiple-input multiple-output (mMIMO), carrier aggregation (CA) and joint mMIMO and CA capabilities, respectively.
In [21], the authors compared the energy efficiency trade-offs associated with a macrocellular network equipped with shutdown capabilities to those of a small cell network, discussing the pros and cons of these two deployment strategies The further analysis in [22] showed the benefits of deploying advanced sleep modes at the small cells, proposing multiple carrier shutdown strategies, and comparing them in terms of complexity, blocking rate probability, throughput and energy efficiency.
From an optimization perspective, the authors of [23] investigated how to fine-tune network energy efficiency by jointly managing the long-term cell activation, UE association and power control in a heterogeneous network with mMIMO capabilities, and proposed a distributed solution based on game theory, proven to converge to the Nash equilibrium.
The work in [24] also modelled hardware failure rate due to cell (de)activation, and proposed a heuristic to control the BS shutdown, which minimises the acceleration over time of the hardware failure rate, while satisfying the UE quality of service demands.
Although the models produced and insights gained from these studies are valuable, their reliance on numerous assumptions —stemming from the theoretical nature and complexity of large-scale networks—limits their applicability. For instance, the carrier shutdown scenarios explored in this research significantly differ from those implemented in actual network environments. Therefore, these theoretical approaches fall short when applied to the optimization of energy saving solutions in real-world network settings.
I-A4 Data-driven modelling and optimization
Given the limitations noted, there is an increasing interest in leveraging ML to develop more scalable and flexible approaches for network optimization [25]. Recognizing this potential, the 3GPP has initiated a shift towards a new paradigm in data-driven network modeling and optimization. This effort focuses on identifying the crucial measurements and protocols necessary to devise advanced data-driven strategies for localized network optimization [26]. By applying ML techniques, the aim is to predict network performance accurately and improve various dimensions of network functionality, notably energy efficiency [27].
To date, such explorations into data-driven network modeling and optimization have predominantly focused on developing ML models that harness data from numerical simulations and test environments. These models are not just academic exercises, but are intended to introduce new mechanisms at the physical layer of communication systems [28], as well as to refine the functionalities at the network’s upper layers [29]. Significant advancements in physical layer mechanisms include the design of neural receivers specifically for 5G multi-UE multiple-input multiple-output (MIMO) configurations [30], alongside the introduction of deep learning-based synchronization [31] and decoding schemes [32]. Intelligent MIMO channel estimators have also emerged as a notable example application. These ML estimators are designed to mitigate the complexity inherent in channel detection procedures while minimizing performance degradation [33]. In terms of radio resource management functionalities, new deep learning approaches aim at streamlining wireless resource allocation algorithms, showcasing a significant leap in managing the complexities of medium access control (MAC) layer network operations [34].
In contrast, only a few studies have explored the use of real-world network data for modeling and enhancing large-scale networks, the subject at hand.
Traffic prediction
Within this domain, substantial strides have been made in data-driven traffic prediction with real-world data employing various ML techniques, such as long short term memory cells, convolutional neural networks, and graph convolutional networks [35, 36, 37, 38]. These methodologies are pivotal for facilitating proactive network optimization efforts. Cutting-edge traffic prediction approaches are increasingly leveraging generative adversarial networks, combined with traffic data and expert knowledge [39, 40]. Integration of urban environment knowledge graphs have further enriched traffic prediction models, improving their accuracy and thus the quality of network optimization.
Anomaly detection
The application of data-driven modeling using real-network data has also been significantly explored for anomaly detection and network alarm predictions. Advanced ML models, similar to those used for traffic prediction, have been adapted to detect anomalies by learning from vast datasets of network activity. These models are trained to recognize patterns indicative of potential issues, enabling network operators to preemptively address problems before they escalate [41, 42, 43, 44].
Cell rate and UE spectral efficiency predictions
Gijon et al. assessed cell throughput prediction accuracy using real network key performance indicator (KPI) counters, exploring a variety of ML methodologies such as support vector regression, k-nearest neighbors, and decision trees and artificial neural network (ANN)-based ensemble approaches [45]. Advancing this research, Xing et al. introduced a neural network framework aimed at modeling the interference and spectral efficiency of downlink (DL) 5G mMIMO transmissions, employing reference signal received power (RSRP) data from DTs [46]. Their work underscores the potential of ML in accurately forecasting cell throughput and enhancing understanding of UE spectral efficiencies. However, these models lack the capability to dynamically forecast KPIs in response to changes in network configurations, revealing a significant gap in their adaptability to real-world optimization challenges. Furthermore, they face challenges in accurately predicting UE rates, mainly because they do not adequately consider crucial factors such as bandwidth allocation per UE, which are profoundly influenced by network energy-saving strategies.
Energy consumption and carbon emissions estimations
To evaluate the energy-related performance of the RAN, the mobile industry has defined several measurement methods and metrics for different network levels (network [47], site [48], BS [49], and UE [50]) as well for different scenarios (dense urban, urban, and rural coverage) [47], and services (enhanced mobile broadband, ultra-reliable low-latency communications, and massive machine type communications) [51]. More recently, Rappaport et al. have introduced the Waste Factor [52], a new metric for quantifying energy efficiency in a wide range of circuits and systems applications, including data centers and RANs.
Li et al.’s study, utilizing network data from Nanchang, modeled carbon emissions from mobile networks and expanded these insights across China [53]. The research used energy consumption and network traffic data from Nanchang, alongside BS and mobile user counts across Chinese provinces, using simple models for extrapolation. Additionally, they introduced a reinforcement learning (RL) framework designed to shut down network cells as a means to prevent capacity over-provisioning. A significant limitation of the model, however, is its simplistic approach to energy-saving schemes —cells are shut down as deemed necessary and UE handovers are ideal— coupled with an inability to evaluate UE performance following changes. Maggi et al. have made a substantial contribution by introducing a sector-specific, closed-loop Bayesian method to optimize carrier shutdown thresholds, effectively reducing power consumption while maintaining network KPIs, particularly UE rates [54]. This data-driven approach, validated in a live 4G network, emphasizes the practical application and effectiveness of their model. However, the focus on optimizing carrier shutdown thresholds, excluding the refinement of handover algorithms, and the model’s limitation in simulating network behavior offline might present challenges in more complex scenarios where Bayesian optimization may fail to converge [55].
Drawing from the literature, it becomes evident that while significant strides have been made in network optimization through various methodologies, a crucial gap remains: a cohesive, data-driven modeling framework. Existing research, valuable though it is,often compartmentalize aspects of network performance, neglecting the intricate balance required between energy efficiency and maintaining optimal network KPIs in dynamic real-world scenarios. This highlights the urgent need for a comprehensive framework that not only bridges these divides, but also utilizes real-world data to thoroughly model and optimize the trade-offs between energy consumption and key network performance metrics. This includes considering the complex interplay between energy-saving strategies, mobility management, and other network dynamics. Such a framework promises to advance the current state of network optimization by providing a more accurate, adaptable, and practical approach to energy-saving solutions, thereby addressing the limitations identified in the existing body of work.
I-B Research questions and objectives
In light of the current state of the art, this paper sets forth a series of research questions and scientific objectives to delineate the research problem at hand, and move towards the creation of such comprehensive end-to-end framework:
Research Questions:
  • Is it feasible for a data-driven model, reliant solely on off-the-shelf network data —no DTs— to deliver accurate predictions of network energy consumption and user throughput?
  • How can we seamlessly blend machine learning algorithms with domain expertise to enhance energy efficiency modeling in large-scale networks?
  • What advancements in network energy efficiency prediction can this type of modeling framework introduce, and how does it measure up against existing methodologies?
Scientific objectives:
  • Establish a Data-Driven Modeling Framework: Our goal is to devise a holistic framework that can precisely forecast network energy efficiency and user throughput using only inexpensive available network data. This requires the harmonization of machine learning (ML) technologies with expert insights to depict the multifaceted nature of large-scale networks.
  • Consider Practical Network Energy Saving Solutions: The framework aims to furnish practical energy saving solutions into the large-scale network modelling for elevating energy efficiency in telecom networks.
  • Validate the Framework with Real-World Data: We intend to prove the framework’s utility by applying it to actual network scenarios, showcasing its capability for precise energy consumption predictions and its value in fostering user throughput when fine tuning energy saving solutions.
I-C Paper’s contributions
@@
Building upon the research questions and scientific objectives outlined previously, this paper introduces a pioneering solution to the challenges of optimizing network energy-saving solutions. We present a novel data-driven framework for modeling network energy consumption and UE throughput, referred to as SRCON [56], which uses measurement data from live networks to jointly fit (i) ML-based, black-box and (ii) expert-based, white-box models222 In contrast to a black-box model, in a white-box model logic, functioning and programming steps are transparent and known. As a result, it’s decision making process is interpretable. that can
@@
  • generalise the modelling of network energy consumption and UE throughput to any possible network energy saving solution configuration, even those not observed in the training data.
SRCON, akin to digital twins, functions as a virtual representation of the physical system; in this instance, the network, acting as its digital counterpart. Digital twins are increasingly viewed as a vital paradigm for monitoring, controlling, and optimizing communication systems [57, 58]. They offer a sandbox for testing new features, e.g. artificial intelligence (AI) solutions, potentially reducing the need for field data collection and algorithm testing. Most digital twins often rely on Internet of things (IoT) sensors to proactively analyze system performance in real-time and make decisions accordingly [59]. A key challenge in digital twin is ensuring that virtual control optimization is safe and reliable, preventing incorrect decisions due to ”model exploitation.” [60] SRCON diverges from typical digital twin frameworks —and attempts to enhance reliability— by not attempting to replicate network behaviors on a small timescale, e.g., milliseconds —a task that is fundamentally impractical due to the significant modeling challenges faced in real-world large-scale networks. Unlike the relatively straightforward sensor data utilized in many digital twin applications, the network’s ’sensors’ are the UEs themselves, whose interactions and performance data are inherently multi-dimensional and complex to analyze. Given this intricacy, SRCON instead is set to achieve a statistically indistinguishable emulation of network behaviors. We should also underscore that, unlike many digital twins focused on real-time optimization, SRCON’s current goal is not real-time modeling and optimization but rather to establish foundational functions for such future optimization capabilities. Further details on these aspects are provided throughout the document.
It is important to acknowledge that our current scope focuses on network energy consumption and UE throughput, pivotal KPIs for evaluating the efficiency of energy-saving solutions. However, we recognize the significance of other metrics, such as delay, Diving into the modeling of such metrics represents a promising avenue for future research, requiring additional modelling. We leave this to future work.
@@
With respect to the available literature earlier reviewed in Section I-A, the novelty of this paper resides in the following aspects:
@@
  • Novel Modeling Approach: Unlike existing literature, this work uniquely focuses on modeling network energy consumption and UE throughput within the context of a practical carrier shutdown solution and its associated handover procedures across a vast, heterogeneous network of hundreds of 4G and 5G cells. This eliminates the need for expensive drive testing or ray tracing-based simulation tools.
@@
  • Decomposed Modeling Framework: We introduce a novel methodology by segmenting the overall modeling challenge into distinct ML- and expert-based subproblems. Outputs from expert-based models inform ML-based models, enabling precise generalization of network energy consumption and UE throughput predictions across any carrier shutdown configuration, including unencountered scenarios.
  • Data and Model Construction Pipeline: This paper provides a detailed account of the necessary data and the pipeline used to construct ML-based models for network energy consumption and UE throughput. This method is crucial for accurately capturing the varied performance characteristics of network products and end-user devices — an approach not seen in current literature. Each problem necessitates a unique framework.
  • Minimal Assumption Requirement: A distinguishing feature of our modeling approach is its minimal dependence on the statistical assumptions often mandated by methods like Bayesian optimization, which typically assume Gaussian behavior. Our models’ capability to provide reliable predictions without such assumptions significantly increases their robustness and adaptability across various network scenarios.
  • Innovative Expert-Based ABM Model: Introducing an innovative agent-based model (ABM) marks a significant advancement. This expert model thoroughly simulates network stochasticity and the dynamics of carrier shutdown across wide-area networks with manageable complexity. It includes handover procedures and approximates UE transfers. The ABM enables precise predictions using established telecom theories, supplying inputs to ML models for any configuration of carrier shutdown solutions and handover parameters.
@@
  • Data-Driven Comparative Results: We conclude with empirical findings that underscore the extensive benefits of our modeling approach. A comparative evaluation against a benchmark tool, utilized by a leading network operator for carrier shutdown optimization, illustrates our models’ superior performance and effectiveness.
@@ II-B Reference carrier shutdown (de)activation logic
In the reference implementation used in this paper, assuming that a capacity cell, $c$, is paired with only one coverage cell, $b(c)$, and to maintain the quality of service (QoS) of UEs affected by the transfer, carrier shutdown can only be autonomously activated in capacity cell, $c$, when
@@
The careful consideration of these algorithms ensures that the paired coverage cells can seamlessly support the communications of all UEs in the capacity cell that is shutting down, in addition to those of its already connected UEs, thereby maintaining QoS without disruption. A bias could be added to the corresponding thresholds to ensure that a larger number of PRBs are reserved for the UEs being transferred. By ensuring that the paired coverage cells can accommodate the PRBs of the transferred UEs, we uphold the QoS.
@@ III Network Energy Efficiency Modelling
As can be inferred from previous sections, the network energy efficiency modelling problem deals with multiple objectives (e.g. network energy consumption and UE throughput) and optimization variables (e.g. carrier shutdown and A4 handover parameters), and is particularly challenging due to its large-scale, stochasticity and non-stationarity as well as the complex coupling between cells and the intricate trade-offs between energy consumption and UE performance. For instance, the shutdown of a cell may save energy, but impacts, not only the performance of the UEs connected to such cell, but also the overall network coverage and the performance of those UEs connected to the neighboring cells as well as the possibility of nearby capacity cells to shutdown. In practice, these challenges make the utilization of precise networking models essential to perform a rigorous network energy efficiency optimization.
@@
To address these challenges, SRCON —our proposed modelling approach— combines a variety of wireless as well as data-driven and ML concepts. Rather than attempting to replicate network behaviours at a time-scale of milliseconds, which is fundamentally infeasible, considering the aforementioned modelling challenges in practical networks, SRCON emulates network behaviours in a statistically indistinguishable manner. This statistical indistinguishability lays the ground for a practical and accurate overall assessment of UE performance, allowing the use of sufficient statistics —instead of all available samples— through an efficient processing of massive data sets. To enable an accurate generalization, and deal with the heterogeneity and complexity of devices, SRCON uses measurement data from live networks to jointly fit ML-based, black-box and expert-based, white-box models, as hinted earlier. For our network energy efficiency modelling problem, in more details, we propose to model BS energy consumption and UE throughput statistics using customised ML-based modeling approaches to the available data to capture the particularities of the specific BS products and off-the-shelf UEs in the area of study (e.g. not all BS product versions have the same energy consumption characteristics, not all UEs have the same decoding capabilities). A systematic feature importance analysis was used to identify the most relevant input features of such two ML-based models, e.g. the amount of time for which capacity cells will be in carrier shutdown, the number of UEs transferred to each neighbouring cell when carrier shutdown takes place, the resulting PRB load in those neighbouring cells. To allow generalisation, a customised expert-based model is used to derive the inputs of the ML-based models for any carrier shutdown and A4 handover parameter configuration. In other words, said white-box model takes as input, among others, the carrier shutdown and A4 handover parameter configuration, and provides as output the inputs of our black-box models. In this manner, we can drive universal ML-based BS energy consumption and UE throughput predictions, and further derive network energy efficiency.